The Nut Barn needs 10 pounds of mixed peanuts and cashews to sell for $8.50 per pound. Cashews alone sell for $10 per pound and peanuts for $7.50 per pound. What is the amount of cashews that can be in the mixture and hold the price to $8.50 per pound?
a. Choose a variable to represent the number of pounds of peanuts. Then find the number of pounds of cashews to be added to make 10 pounds total. Now figure out the cost of the amount of each nut in the mixture.
b. Write the equation.
c. Solve the equation

Answer by checkley71(8403)

10x+7.50(10-x)=8.50*10
10x+75-7.50x=85
2.5x=85-75
2.5x=10
x=10/2.5
x=4 pounds of cashews are needed.
10*4+7.5(10-4)=85
40+7.5(6)=85
40+45=85
85=85